# Reconstruction of strong degeneracy region for parabolic equations and systems

### v1

@@ Abstract
We address the inverse problem of recovering a degeneracy point within the diffusion coefficient of a one-dimensional complex parabolic equation by observing the normal derivative at one point of the boundary. The strongly degenerate case is analyzed. In particular, we derive sufficient conditions on the initial data that guarantee the stability and uniqueness of the solution obtained from a one-point measurement. Moreover, we present more general uniqueness theorems, which also cover the identification of the initial data and the coefficient of the zero order term, using measurements taken over time. Our method is based on a careful analysis of the spectral problem and relies on an explicit form of the solution in terms of Bessel functions. Our investigation also covers the case of real 1-D degenerate parabolic systems of equations coupled with a specific structure. Theoretical results are also supported by numerical simulations.
Keywords: inverse problems, degenerate parabolic equations, numerical reconstruction.
@@ 1 Introduction
For example, the inverse problem of reconstructing an interior degeneracy point for the real case was considered in [5], where the authors analyzed the strongly degenerate case for $\theta=1$. Our goal is to generalize this result in two directions. One direction involves extending it to systems of two real degenerate coupled parabolic equations. Specifically, in this work, we consider a coupling with a particular structure that allows us to reformulate the problem as a complex degenerate parabolic equation. The other direction concerns extending the result to cases where the degeneracy has an exponent different from 1. In particular, we will analyze the strongly degenerate case with $\theta\in[1,2)$. Here, the analysis is technically more complicated and our results generalize those in [5]. Instead, the weakly degenerate case with $\theta\in[0,1)$ is still an open problem. In fact, we cannot analyze the spectral behavior independently to the left and right of the degeneracy. This distinction prevents a spectral analysis analogous to what is possible in the strongly degenerate situation, with the exception of specific configurations of the degeneracy point. Furthermore, from an applicative point of view, as in climatology or financial mathematics, the most interesting examples fall into the class of strongly degenerate problems.
It should also be noted that the restriction $\theta<2$ is due to the spectral technique implemented, specifically to the use of Bessel functions.
@@
The paper is organized as follows. In Section 2, we introduce the functional setting and establish the well-posedness of the corresponding direct problem. In Section 3, we analyze the eigenvalue problem. In Section 4, we provide an expression of the normal derivative computed using Bessel functions. In Section 5, we establish a Lipschitz stability result with one-point measurements. Section 6 is devoted to general uniqueness results for distributed measurements over a time interval, assuming that the initial data and the coefficient $c$ are also unknown. Section 7 concerns the application of the previous results to real systems of 1-D coupled degenerate parabolic equations. Section 8 concludes with numerical experiments related to the inverse problem under consideration.
@@ 2 Functional setting and well-posedness
Proof of Proposition 2.2: The proof is analogous to that in the real case (see, for instance, [8] and [9]). $\blacksquare$
@@ 3 The eigenvalue problem
The analysis of the spectral problem associated to (1) will be essential for our purposes. The eigenvalues and associated eigenfunctions of the degenerate diffusion operator $w\mapsto-(|x-a|^{\theta}w^{\prime})^{\prime}$ are the nontrivial solutions $(\lambda,\phi)$ of
@@
by means of the two coordinate transformations $y=\dfrac{x-a}{1-a}$, with $\varphi(y)=\phi(a+(1-a)y)$, and $y=\dfrac{x-a}{a}$, with $\varphi(y)=\phi(a+ay)$, respectively. The first eigenvalue sub-problem can be rewritten as a differential Bessel’s equation of order $\nu_{\theta}=\frac{\theta-1}{2-\theta}$
@@ 4 Computation of the normal derivative
$\begin{cases}-((x-a)^{\theta}\phi^{\prime}(x))^{\prime}=\lambda\phi,\quad x\in(a,1),\\[2.84526pt] \phi(1)=0,\quad(x-a)^{\theta}\phi^{\prime}(x)\big{|}_{x=a}=0\end{cases}$
@@ 5 Lipschitz stability with one point measurement
$=\Bigg{|}e^{\alpha t}R(\beta t)\Bigg{(}\dfrac{2k_{\theta}e^{-\lambda_{1}t}}{d_{\nu_{\theta},1}}\begin{pmatrix}F^{1}_{1}(a)\\ F^{2}_{1}(a)\end{pmatrix}+\sum_{n=2}^{\infty}\dfrac{2k_{\theta}e^{-\lambda_{n}t}}{d_{\nu_{\theta},n}}\begin{pmatrix}F^{1}_{n}(a)\\ F^{2}_{n}(a)\end{pmatrix}\Bigg{)}\Bigg{|}\$,
@@ 6 Uniqueness results for “distributed”measurements
In this section, we present sufficient conditions for the uniqueness result of two more general inverse problems. In the first one, we consider the degeneracy point, initial data, and coefficient $c$ as unknowns; in the second one, only the degeneracy point and the initial data are unknown. Unlike in the previous section, where we considered point-wise measurements of $\partial_{x}w(1,t)$, here we require measurements distributed over a time interval. For the first inverse problem, in addition to distributed measurements of $\partial_{x}w(1,t)$, the uniqueness of the coefficient $c$ also requires distributed measurements of $\partial_{x}w(0,t)$, which also allow to achieve the uniqueness of the initial data over the entire interval $(0,1)$. The second inverse problem simply requires distributed measurements of $\partial_{x}w(1,t)$, but the uniqueness of the initial data from the $\partial_{x}w(1,t)$ measurements is confined to the right subinterval $(a,1)$.
@@
where now also the initial datum $w_{0}$ and the coefficient $c=\alpha+i\beta$ are unknowns. The solution and the normal derivative for the left sub-problem (37) can be computed with an analysis similar to that of Section 4, obtaining
@@ 7 Real systems of 1-D coupled degenerate parabolic equations
where
$\lambda_{1}(\gamma)=k_{\theta}^{2}\dfrac{j_{\nu_{\theta},1}^{2}}{(1-\gamma)^{2k_{\theta}}}\ .$
@@
The normal derivatives at the boundary $\begin{pmatrix}\partial_{x}u(0,t)\\ \partial_{x}v(0,t)\end{pmatrix}$ are determined using (40). Once again, we can distinguish between two uniqueness Theorems.

### v2

# Reconstruction of degeneracy region and power for parabolic equations and systems

@@ Abstract
We address the inverse problem of recovering a degeneracy point within the diffusion coefficient of a one-dimensional complex parabolic equation by observing the normal derivative at one point of the boundary. The strongly degenerate case is analyzed. In particular, we derive sufficient conditions on the initial data that guarantee the stability and uniqueness of the solution obtained from a one-point measurement. Moreover, we present more general uniqueness theorems, which also cover the identification of the initial data, the coefficient of the zero order term and the degeneracy power, using measurements taken over time. Our method is based on a careful analysis of the spectral problem and relies on an explicit form of the solution in terms of Bessel functions. Our investigation also covers the case of real 1-D degenerate parabolic systems of equations coupled with a specific structure. Theoretical results are also supported by numerical simulations.
footnotetext: Keywords: inverse problems, degenerate parabolic equations, numerical reconstruction.footnotetext: 2020 Mathematics Subject Classification: 35K65, 35R30, 80A23, 65M32.
@@ 1 Introduction
For example, the inverse problem of reconstructing an interior degeneracy point for the real case was considered in [7], where the authors analyzed the strongly degenerate case for $\theta=1$. Our goal is to generalize this result in several directions. One direction involves extending it to systems of two real degenerate coupled parabolic equations. Specifically, in this work, we consider a coupling with a particular structure that allows us to reformulate the problem as a complex degenerate parabolic equation. Another direction of generalization concerns extending the result to cases where the degeneracy has an exponent different from 1. In particular, we will analyze the strongly degenerate case with $\theta\in[1,2)$. Here, the analysis is technically more complicated and our results generalize those in [7]. Instead, the weakly degenerate case with $\theta\in[0,1)$ is still an open problem. In fact, we cannot analyze the spectral behavior independently to the left and right of the degeneracy. This distinction prevents a spectral analysis analogous to what is possible in the strongly degenerate situation, with the exception of specific configurations of the degeneracy point. Furthermore, from an applicative point of view, as in climatology or financial mathematics, the most interesting examples fall into the class of strongly degenerate problems. It should also be noted that the restriction $\theta<2$ is due to the spectral technique implemented, specifically to the use of Bessel functions.
The last extension regards the uniqueness and reconstruction for the inverse problems where also the degeneracy power, the coefficient of the zero-order term, and the initial data are unknown. In this case, our result for the degeneracy power also differs from the one in [6], in terms of the measurements considered and for the extension to systems. In particular, in [6] the authors take into account measurements of the time derivative and space gradient for a fixed time, but distributed on the entire spatial domain. Our approach considers a more natural observation, represented by the normal derivative at one boundary distributed over a time interval.
@@
The paper is organized as follows. In Section 2, we introduce the functional setting and establish the well-posedness of the corresponding direct problem. In Section 3, we analyze the eigenvalue problem. In Section 4, we provide an expression of the normal derivative computed using Bessel functions. In Section 5, we establish a Lipschitz stability result with one-point measurements. Section 6 is devoted to general uniqueness results for distributed measurements over a time interval, assuming that the initial data, the coefficient $c$ and the degeneracy power are also unknown. Section 7 concerns the application of the previous results to real systems of 1-D coupled degenerate parabolic equations. Section 8 concludes with numerical experiments related to the inverse problem under consideration.
@@ 2 Functional setting and well-posedness
Proof of Proposition 2.2: The proof is analogous to that in the real case (see, for instance, [10]). $\blacksquare$
@@ 3 The eigenvalue problem
The analysis of the spectral problem associated to (1) will be essential for our purposes. The eigenvalues and associated eigenfunctions of the degenerate diffusion operator A-A are the nontrivial solutions $(\lambda,\phi)$ of
@@
by means of the two coordinate transformations $y=\dfrac{x-a}{1-a}$, with $\varphi(y)=\phi(a+(1-a)y)$, and $y=\dfrac{x-a}{a}$, with $\varphi(y)=\phi(a+ay)$, respectively.
The first eigenvalue sub-problem can be rewritten as a differential Bessel’s equation of order $\nu_{\theta}=\frac{\theta-1}{2-\theta}$
@@ 4 Computation of the normal derivative
$\begin{cases}-((x-a)^{\theta}\phi^{\prime}(x))^{\prime}=\lambda\phi,\quad x\in(a,1),\\[2.84526pt] \phi(1)=0,\quad(x-a)^{\theta}\phi^{\prime}(x)\big|_{x=a}=0\end{cases}$ (8)
@@ 5 Lipschitz stability with one point measurement
[…]
@@ 6 Uniqueness results for “distributed”measurements
In this section, we present sufficient conditions for the uniqueness result of three more general inverse problems. In the first one, we consider the degeneracy point, initial data, and coefficient $c$ as unknowns; in the second one, only the degeneracy point and the initial data are unknown. Moreover, the third inverse problem concerns the case where the unknowns are represented by the degeneracy power and the initial data. Unlike in the previous section, where we considered point-wise measurements of $\partial_{x}w(1,t)$, here we require measurements distributed over a time interval. For the first inverse problem, in addition to distributed measurements of $\partial_{x}w(1,t)$, the uniqueness of the coefficient $c$ also requires distributed measurements of $\partial_{x}w(0,t)$, which also allow us to achieve the uniqueness of the initial data over the entire interval $(0,1)$. The second and third inverse problems simply require distributed measurements of $\partial_{x}w(1,t)$, but the uniqueness of the initial data from the $\partial_{x}w(1,t)$ measurements is confined to the right sub-interval $(a,1)$.
@@
where we take into account different unknowns, depending on the inverse problem we are analyzing. The solution and the normal derivative for the left sub-problem (38) can be computed with an analysis similar to that of Section 4, obtaining
@@
We now state the uniqueness result for the third inverse problem, where the degeneracy power and the initial data are unknown.
Theorem 6.3
Let $\theta_{1},\theta_{2}\in[1,2)$ and $0<t_{1}<t_{2}$. Let $w^{\theta_{1}}$ and $w^{\theta_{2}}$ be two solutions to (37), corresponding to the initial values $w_{0}=u_{0}+iv_{0}$ and $\widetilde{w}_{0}=\widetilde{u}_{0}+i\widetilde{v}_{0}$, respectively. Assume that
$\begin{pmatrix}U_{1}^{0}(a;\theta_{1})\\ V_{1}^{0}(a;\theta_{1})\end{pmatrix}\neq\begin{pmatrix}0\\ 0\end{pmatrix}\qquad\text{and}\qquad\begin{pmatrix}\widetilde{U}_{1}^{0}(a;\theta_{2})\\ \widetilde{V}_{1}^{0}(a;\theta_{2})\end{pmatrix}\neq\begin{pmatrix}0\\ 0\end{pmatrix}$, (45)
where $U^{0}_{1}(a;\theta_{1})$ and $\widetilde{U}^{0}_{1}(a;\theta_{2})$, for $u_{0}$ and $\widetilde{u}_{0}$ respectively, are given by (14) (with $n=1$) and $V^{0}_{1}(a;\theta_{1})$ and $\widetilde{V}^{0}_{1}(a;\theta_{2})$, for $v_{0}$ and $\widetilde{v}_{0}$ respectively, are given by (15) (with $n=1$).
Then $\partial_{x}w^{\theta_{1}}(1,t)=\partial_{x}w^{\theta_{2}}(1,t)$ for $t_{1}<t<t_{2}$ implies that $\theta_{1}=\theta_{2}$ and $u_{0}=\widetilde{u}_{0}$, $v_{0}=\widetilde{v}_{0}$ in $(a,1)$.
Before we prove this uniqueness result, we need to prove the monotonicity of the first eigenvalue with respect to the degeneracy power $\theta$.
6.1 Monotonicity of the first eigenvalue with respect to $\theta$
In this subsection, we will consider the right eigenvalue problem (8) and prove the monotonicity of the first eigenvalue $\lambda_{1}$ w.r.t. $\theta$. Since $\theta<2$, the injection of $H^{1}_{\theta}(a,1)$ into $L^{2}(a,1)$ is compact (see [2] and [8] for the degeneracy coefficient $x^{\theta}$ in $(0,1)$), therefore, we can apply the spectral theorem in the right interval $(a,1)$ and give the variational formulation of the first eigenvalue:
$\lambda_{1}(\theta):=\min_{\phi_{1}\in H^{1}_{\theta}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta}|\phi^{\prime}_{1}|^{2}\,dx}{\int_{a}^{1}|\phi_{1}|^{2}\,dx},$
where $\phi_{1}$ is an associated eigenfunction.
First of all, we want to prove that if $\theta_{1}\leq\theta_{2}$, then $\lambda_{1}(\theta_{1})\geq\lambda_{1}(\theta_{2})$. Observe that, for $x\in(a,1)$, $\theta_{1}\leq\theta_{2}$ implies $(x-a)^{\theta_{2}}\leq(x-a)^{\theta_{1}}$. Hence, taking $\phi_{1}\in H^{1}_{\theta_{1}}(a,1)$, we get
$\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx\leq\int_{a}^{1}\left(x-a\right)^{\theta_{1}}|\phi^{\prime}_{1}|^{2}\,dx$
and
$\|\phi_{1}\|^{2}_{H^{1}_{\theta_{2}}}=\|\phi_{1}\|^{2}_{L^{2}}+\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx\leq\|\phi_{1}\|^{2}_{L^{2}}+\int_{a}^{1}\left(x-a\right)^{\theta_{1}}|\phi^{\prime}_{1}|^{2}\,dx=\|\phi_{1}\|^{2}_{H^{1}_{\theta_{1}}}<+\infty,$
obtaining the continuous injection $H^{1}_{\theta_{1}}\subseteq H^{1}_{\theta_{2}}$. Thus, taking the minimum over $\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}$, we have
$\min_{\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}dx}{\int_{a}^{1}|\phi_{1}|^{2}dx}\leq\min_{\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{1}}|\phi^{\prime}_{1}|^{2}dx}{\int_{a}^{1}|\phi_{1}|^{2}dx}:=\lambda_{1}(\theta_{1}).$
We now want to prove that
$\lambda_{1}(\theta_{2}):=\min_{\phi_{1}\in H^{1}_{\theta_{2}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx}{\int_{a}^{1}|\phi_{1}|^{2}\,dx}=\min_{\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx}{\int_{a}^{1}|\phi_{1}|^{2}\,dx},$ (46)
observing that, due to the injection $H^{1}_{\theta_{1}}\subseteq H^{1}_{\theta_{2}}$, we have
$\lambda_{1}(\theta_{2})\leq\min_{\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx}{\int_{a}^{1}|\phi_{1}|^{2}\,dx}.$ (47)
To obtain the equality, we exploit the density of $H^{1}_{\theta_{1}}$ in $H^{1}_{\theta_{2}}$. In fact, the functions of class $C^{\infty}$ compactly supported in $(a,1)$ are dense in each $H^{1}_{\theta}$ (see [5]). Therefore, from $C^{\infty}_{0}(a,1)\subseteq H^{1}_{\theta_{1}}(a,1)\subseteq H^{1}_{\theta_{2}}(a,1)$, we deduce $\overline{H^{1}_{\theta_{1}}}^{\|\cdot\|_{H^{1}_{\theta_{2}}}}=H^{1}_{\theta_{2}}$. Using this density result, we can also conclude that
$\lambda_{1}(\theta_{2})\geq\min_{\phi_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}}\dfrac{\int_{a}^{1}\left(x-a\right)^{\theta_{2}}|\phi^{\prime}_{1}|^{2}\,dx}{\int_{a}^{1}|\phi_{1}|^{2}\,dx}$
and, using also (47), we deduce the validity of equality (46), from which we get $\lambda_{1}(\theta_{2})\leq\lambda_{1}(\theta_{1})$.
Finally, we want to prove the strict monotonicity of $\lambda_{1}$ w.r.t. $\theta$. Taking into account a minimum point $\tilde{\phi}_{1}\in H^{1}_{\theta_{1}}\setminus\{0\}$ and $\theta_{1}<\theta_{2}$, we obtain $(x-a)^{\theta_{2}}<(x-a)^{\theta_{1}}$ for $x\in(a,1)$. Hence, $\int_{a}^{1}\left(\left(x-a\right)^{\theta_{1}}-\left(x-a\right)^{\theta_{2}}\right)|\tilde{\phi}^{\prime}_{1}|^{2}dx=0$ would imply $|\tilde{\phi}^{\prime}_{1}|^{2}=0$ a.e. on $(a,1)$, which is a contradiction because it would also require $\tilde{\phi}_{1}=0$. Thus, we obtain $\lambda_{1}(\theta_{2})<\lambda_{1}(\theta_{1})$.
Proof of Theorem 6.3: Let us consider the right sub-problem (37) and set
$\lambda_{n}:=k_{\theta_{1}}^{2}\frac{j_{\nu_{\theta_{1}},n}^{2}}{(1-a)^{2k_{\theta_{1}}}},\quad\mu_{n}:=k_{\theta_{2}}^{2}\frac{j_{\nu_{\theta_{2}},n}^{2}}{(1-a)^{2k_{\theta_{2}}}},\quad n\in\mathbb{N},$
for which we have $\lambda_{1}<\lambda_{2}<\ldots$ and $\mu_{1}<\mu_{2}<\ldots$. Assume that $\theta_{1}\neq\theta_{2}$. Without loss of generality, we can assume that $\theta_{1}<\theta_{2}$, then $\lambda_{1}>\mu_{1}$, due to the above monotonicity result. Using, for example, the condition $\partial_{x}w^{\theta_{1}}(1,t)=\partial_{x}w^{\theta_{2}}(1,t)$ and the analyticity w.r.t. time, we now prove that $\theta_{1}$$\theta_{2}$. Therefore, setting
$W^{0}_{n}(a;\theta_{1}):=\begin{pmatrix}U_{n}^{0}(a;\theta_{1})\\ V_{n}^{0}(a;\theta_{1})\end{pmatrix}\quad\text{and}\quad\widetilde{W}^{0}_{n}(a;\theta_{2}):=\begin{pmatrix}\widetilde{U}^{0}_{n}(a;\theta_{2})\\ \widetilde{V}^{0}_{n}(a;\theta_{2})\end{pmatrix}$,
we obtain
$\sum_{n=1}^{\infty}\dfrac{e^{-\lambda_{n}t}}{d_{\nu_{\theta_{1}},n}}R(\beta t)W^{0}_{n}(a;\theta_{1})=\displaystyle\sum_{n=1}^{\infty}\dfrac{e^{-\mu_{n}t}}{d_{\nu_{\theta_{2}},n}}R(\tilde{\beta}t)\widetilde{W}^{0}_{n}(a;\theta_{2})$
for $t>t_{1}$. Therefore, taking the absolute value, we have
$\left|\dfrac{e^{-(\lambda_{1}-\mu_{1})t}}{d_{\nu_{\theta_{1}},1}}W^{0}_{1}(a;\theta_{1})+\displaystyle\sum_{n=2}^{\infty}\dfrac{e^{-(\lambda_{n}-\mu_{1})t}}{d_{\nu_{\theta_{1}},n}}W^{0}_{n}(a;\theta_{1})\right|=\left|\dfrac{1}{d_{\nu_{\theta_{2}},1}}\widetilde{W}^{0}_{1}(a;\theta_{2})+\displaystyle\sum_{n=2}^{\infty}\dfrac{e^{-(\mu_{n}-\mu_{1})t}}{d_{\nu_{\theta_{2}},n}}\widetilde{W}^{0}_{n}(a;\theta_{2})\right|.$
Since $\lambda_{1}>\mu_{1}$, we let $t\to\infty$ in the previous equality to get $\left|\widetilde{W}^{0}_{1}(a;\theta_{2})\right|=0$, in contrast with (45). Thus, $\theta_{1}=\theta_{2}$ follows, and this implies $\lambda_{n}=\mu_{n}$ for all $n\in\mathbb{N}$. The uniqueness of the initial data in $(a,1)$ can be achieved following the same reasoning as in the Proof of Theorem 6.1.
Remark 6.2
We observe that the above method can be adapted to obtain a uniqueness result for a more general class of problems, even in abstract form. For this purpose one could consider a family of differential operators such that the corresponding first eigenvalues satisfy a suitable monotonicity result, w.r.t. the unknown parameter that we are reconstructing, and a non degenerate observation of all first eigenfunctions exists. We will not further develop such an investigation in this paper, leaving it for a forthcoming study.
@@ 7 Real systems of 1-D coupled degenerate parabolic equations
The normal derivatives at the boundary $\begin{pmatrix}\partial_{x}u(0,t)\\ \partial_{x}v(0,t)\end{pmatrix}$ are determined using (41). Once again, we can distinguish between three uniqueness Theorems.
@@
Theorem 7.4
Let $\theta_{1},\theta_{2}\in[1,2)$ and $0<t_{1}<t_{2}$. Let $\begin{pmatrix}u^{\theta_{1}}\\ v^{\theta_{1}}\end{pmatrix}$ and $\begin{pmatrix}u^{\theta_{2}}\\ v^{\theta_{2}}\end{pmatrix}$ be two solutions to (49), corresponding to the initial values $\begin{pmatrix}u_{0}\\ v_{0}\end{pmatrix}$ and $\begin{pmatrix}\widetilde{u}_{0}\\ \widetilde{v}_{0}\end{pmatrix}$, respectively. Assume that
$\begin{pmatrix}U_{1}^{0}(a;\theta_{1})\\ V_{1}^{0}(a;\theta_{1})\end{pmatrix}\neq\begin{pmatrix}0\\ 0\end{pmatrix}\qquad\text{and}\qquad\begin{pmatrix}\widetilde{U}_{1}^{0}(a;\theta_{2})\\ \widetilde{V}_{1}^{0}(a;\theta_{2})\end{pmatrix}\neq\begin{pmatrix}0\\ 0\end{pmatrix}$,
where $U^{0}_{1}(a;\theta_{1})$ and $\widetilde{U}^{0}_{1}(a;\theta_{2})$, for $u_{0}$ and $\widetilde{u}_{0}$, respectively, are given by (14) (with $n=1$); and $V^{0}_{1}(a;\theta_{1})$ and $\widetilde{V}^{0}_{1}(a;\theta_{2})$, for $v_{0}$ and $\widetilde{v}_{0}$, respectively, are given by (15) (with $n=1$). Then $\begin{pmatrix}\partial_{x}u^{\theta_{1}}(1,t)\\ \partial_{x}v^{\theta_{1}}(1,t)\end{pmatrix}=\begin{pmatrix}\partial_{x}u^{\theta_{2}}(1,t)\\ \partial_{x}v^{\theta_{2}}(1,t)\end{pmatrix}$ for $t_{1}<t<t_{2}$ implies that $\theta_{1}=\theta_{2}$ and $u_{0}=\widetilde{u}_{0}$, $v_{0}=\widetilde{v}_{0}$ in $(a,1)$.